What is the expected return and standard deviation for the following stock?

State of

Economy

Probability of

State of Economy

Rate of Return

if State Occurs

Boom

.06

−.06

Normal

.74

.07

Recession

.20

.18

A.

8.42 percent; 5.69 percent

B.

8.42 percent; 5.74 percent

C.

8.53 percent; 5.74 percent

D.

8.53 percent; 5.69 percent

Please write out by hand/type not in excel
Answer #1

Expected return=Respective return*Respective probability

=(0.06*-6)+(0.74*7)+(0.2*18)=8.42%

probability Return probability*(Return-Expected Return)^2
0.06 -6 0.06*(-6-8.42)^2=12.476184
0.74 7 0.74*(7-8.42)^2=1.492136
0.2 18 0.2*(18-8.42)^2=18.35528
Total=32.3236%

Standard deviation=[Total probability*(Return-Expected Return)^2/Total probability]^(1/2)

=5.69%(Approx).

Hence the correct option is A.